# Resummed Distribution Functions: Making Perturbation Theory Positive and Normalized

### v1

@@ Abstract
Fixed-order perturbative calculations for differential cross sections can suffer from non-physical artifacts: they can be non-positive, non-normalizable, and non-finite, none of which occur in experimental measurements. We propose a framework, the Resummed Distribution Function (RDF), that, given a perturbative calculation for an observable to some finite order in $\alpha_{s}$, will “resum” the expression in a way that is guaranteed to match the original expression order-by-order and be positive, normalized, and finite. Moreover, our ansatz parameterizes all possible finite, positive, and normalized completions consistent with the original fixed-order expression, which can include NnLL resummed expressions. The RDF also enables a more direct notion of perturbative uncertainties, as we can directly vary higher-order parameters and treat them as nuisance parameters. We demonstrate the power of the RDF ansatz by matching to thrust to $\mathcal{O}(\alpha_{s}^{3})$ and extracting $\alpha_{s}$ with perturbative uncertainties from fitting the RDF to ALEPH data.
@@ 1 Introduction
However, at any finite order $M$, our physical perturbative predictions for $p(x)$ may violate unitarity by being non-normalizable or negative. Moreover, the perturbative expansion can be spoiled if $p_{m}(x)\sim\frac{1}{\alpha^{m}}$, which is especially the case in quantum chromodynamics (QCD) where large logarithms due to infrared and collinear divergences can arise. Additionally, while the perturbative series will be an integrable distribution, it will rarely be a proper function and contain objects such as $\delta$-functions or ++-functions (which can technically restore unitarity or cancel divergences). Experimental measurements of $p(x|\alpha)$ are typically true functions that are finite everywhere as nature carries out all-orders calculations, and no experimental histogram will contain these objects.
@@
In this paper, we introduce a new ansatz, which we call the Resummed Distribution Function (RDF), for parameterizing proper PDFs of random variables. This ansatz is guaranteed to be positive, finite, and normalized, and thus satisfy unitarity. The RDF (which we denote as $q(x|\alpha)$ where $x$ can be either a single observable or a set of several observables) is an all-orders expansion in $\alpha$, and it effectively “resums” the fixed-order series using only unitarity and analyticity in $\alpha$ as consistency conditions. Given a fixed-order expansion for $p(x|\alpha)$ up to order $M$, our ansatz $q(x|\alpha)$ is guaranteed to match $p(x|\alpha)$ up to order $M$, meaning that it contains all of the information in the corresponding $p_{\rm FO}(x|\alpha)$. Through this ansatz, we effectively parameterize, in terms of free parameters, all possible higher-order extensions of $p(x|\alpha)$ permissible by unitarity. The RDF provides an automated approach for generating these potential all-orders “resummations” without any appeal to the structure of the theory, making it especially useful if a proper resummation is difficult or unavailable.
@@
The RDF provides an efficient machine-learning-inspired [^1] We say “inspired” because while the RDF is a universal probability estimator with parameters that can be numerically fit with gradient descent like many machine learning methods, our aim is to build a primarily analytical intuition for it. framework to learn normalized, finite, and conditional multi-variable PDFs from analytic calculations, simulations, or real data, complementary to established machine-learning-and-adjacent methods for density estimation. We emphasize the strong inductive bias of the RDF ansatz: while it is a universal PDF approximator, it is particularly useful for distributions that admit a perturbative expansion in $\alpha$, as is often the case in a QFT. In contrast to normalizing flows [^2] We strongly considered naming our ansatz the “Re-normalizing flow”. papamakarios2021normalizingflowsprobabilisticmodeling and diffusion models ho2020denoisingdiffusionprobabilisticmodels , our ansatz is extremely lightweight: it can fit to numeric samples in a matter of minutes, requiring only a small set ($\sim\mathcal{O}(10)$) of trainable parameters. The RDF has a manifest expansion in $\alpha$ where terms of the expansion are directly parameterized, unlike normalizing flows and diffusion models where the expansion is not manifest or necessarily well-behaved. Compared to non-parametric methods for PDF estimation such as Kernel Density Estimation kde , our method can easily be made conditional on a parameter $\alpha$.
@@ 2 Resummed Distribution Functions: A Normalized, Positive, and Finite Ansatz for PDFs
Here, the functions $g$ can be any functions that (a) are bounded from above as any $t_{i}\to\infty$ and (b) are either analytic in $\alpha$ or have single-log dependence on $\alpha$. This parameterization ensures that the $f$ functions are always positive and analytic in $\alpha$, which is necessary for our ansatz to be a valid and analytic PDF (which we discuss further in the following section). [^4] Note that the functions $f$ and $g$ are ordinary functions of $t$ and $\alpha$, not PDFs of $t$ conditioned on $\alpha$. In Eq. (3), the $j^{\rm th}$ integral is only integrating over the $t_{j}$ argument of $f_{j}$, and all other arguments are not integrated over. The ansatz is completely specified by a choice of $g$ functions. Any choice of $g$ functions satisfying the above two conditions is a valid PDF, and all PDFs correspond to some choice of $g$.
@@ 2.1 Univariate Taylor-Expandable Probability Distributions
where $g_{mn}$ are coefficients, and $N$ is the highest power of $t$ considered. The only requirement is that $g_{mN}<0$ for normalization so that $g$ remains bounded from above. With this parameterization, the each part $g$ is a series of polynomials in $t$ and $\alpha$, modulated by $\Theta$-functions in cases where $t$ does not necessarily start at zero. [^5] We draw a comparison here to Ref. Assi:2025ibi , which also involves determining parameters in an exponential to match to a calculation. However, Ref. Assi:2025ibi matches to already-resummed moments, and our $g$’s are inside of an exponential-integral-exponential structure rather than a plain single exponential.
@@ 2.2 Multivariate Taylor-Expandable Probability Distributions
The ordering of the $t_{i}$ is arbitrary, and thus this decomposition is not unique. [^6] Alternatively, one can choose to take an appropriately normalized sum over all possible orderings, but this is expensive computationally and doesn’t offer any great advantage in our case.
@@ 2.3 Analytic Matching
The higher-order information in the ansatz will enforce normalization, positivity, and finiteness, effectively performing an “all-orders” resummation. For the following analytic matching procedure, we assume that we have access to an explicit expression for $p_{\rm FO}(t|\alpha)$ and the coefficient functions $p_{m}$ as defined in Eq. (1). Moreover, we require that the $p_{m}$ are indeed the Taylor coefficient functions of a well-defined distribution; that is, there exists an actual all-orders distribution the $p_{m}$ eventually converge to. [^7] This is not the case if, for example, one has $p_{0}(t)=1$ and $p_{m}(t)=0$ for all $m>0$. This does not converge to a valid distribution over $t$. Note that, in actual QFT calculations and especially QCD, this assumption may not be true: not only is there nonperturbative physics, but often perturbative expansions are only asymptotic series that do not actually converge. These issues are not unique to our method and we do not claim to solve either problem — the RDF is only as good as the perturbation theory used to derive $p_{\rm FO}$.
@@
Notice the $\mathcal{O}(\alpha^{M+1-m^{*}})$ inside the parentheses – since we have divided out $\alpha^{m^{*}}$, the expression inside the parentheses is an $\mathcal{O}(\alpha^{M-m^{*}})$ expression. We highlight this as it will be essential in our $\alpha$-counting. Substituting Eq. (16) into Eq. (15), we have
@@ 2.4 Numeric matching
where the first term corresponds to $\log(g^{*}(t,\alpha))$ and the second term to $g_{\rm Analytic}(t,\alpha)$. The numeric RDF differs from the analytic formula of Eq. (11) in a few minor ways. First, we allow the argument of the logarithm to be an arbitrary polynomial in $\alpha$, rather than using only the lowest power. We also force it to be strictly positive using with an absolute value function. [^9] Why absolute value and not anything else, e.g. a smoothed-out ReLU agarap2019deeplearningusingrectified to enforce positivity? We tried, and it was not as numerically stable. Second, we include explicit factorial scaling of the $g_{mn}$ coefficients for numerical stability. Lastly, we “smooth out” the $\Theta$-functions and absolute value functions, as indicated by the subscript $T$. We do this with the replacements $\Theta_{T}(x)\to\sigma(x;T)=\frac{1}{1+e^{-x/T}}$ and $|x|_{T}\to 2\sigma(\frac{x}{T})x-x$, with learnable parameters $T^{*}_{m}$, $T_{{\rm Analytic},m}$ and $T_{{\rm Abs},m}$ respectively per each $m$. This ensures that the derivatives of the RDF are well-defined at all points in $t$-space, which is especially important for the $\theta_{m}$’s to be learnable via gradient descent. All in all, we fit seven different objects: $g^{*}_{mn}$, ${g_{\mathrm{Analytic}}}_{mn}$, $\theta^{*}_{m}$, ${\theta_{\mathrm{Analytic}}}_{m}$, $T^{*}_{m}$, $T_{\mathrm{Analytic},m}$, and $T_{\textrm{Abs},m}$. For convenience, we will refer to this set of seven learnable objects as $\phi$.
@@
The binwise error in the denominator comes from uncertainties on the data or MC to fit. We then backpropagate to calculate gradients of the loss with respect to the parameters and update the $g_{mn}$ and $\theta_{m}$ arrays with a gradient optimizer, in our case Adam adam .
@@ 3 What, When, and Why RDFs?
In this section, we hope to make clear what the RDF can and cannot do. It is important to be careful about the interpretation of the RDF, as it is not a method for obtaining extra information beyond unitarity for “free” or for performing automatic logarithmic resummation. At the lowest level, the only thing a RDF does is answer the question: “Given a fixed-order calculation, plus the additional knowledge that the all-orders distribution should be a proper probability distribution, what are all probability distributions consistent with that calculation?” The non-matched part of $g$ is infinitely flexible, and so without any additional information, further calculation, or a Bayesian prior, one cannot pick out which of these distributions is the correct one.
@@
However, there is some information gain due to the unitarity assumption: unitarity is an all-orders statement, so the RDF effectively provides a constraint on what the higher-order terms could be. This is indeed a nontrivial constraint: for example, after matching with a RDF to order $M$, and Taylor-expanding to see how the higher order $q_{m}$ for $m>M$ depend on the choice of $g$, one will never find that $q_{m}=0$ for all m>Mm>M [^10] Except in the trivial case where the original fixed-order expression happened to have no $\alpha$-dependence and was normalized to begin with., regardless of the choice of $g$, since that would violate unitarity. No matter what the original $p_{m}$ are, the higher-order coefficients must always conspire to make the entire distribution positive and normalized, which a generic all-orders guess may not do.
The RDF is an expansion in $\alpha^{m-m^{*}}\frac{p_{m}(t)}{p_{m^{*}}(t)}$, where $p_{m}(t)$ is the order-$m$ term in the $\alpha$ expansion and $p_{m^{*}}$ is the first nontrivial term in the expansion. Specifically, we assume in Eq. (18) that $\alpha^{m-m^{*}}\frac{p_{m}(t)}{p_{m^{*}}(t)}\ll 1$ in expanding the logarithms. This is the “resummation” in “Resummed Distribution Function”: it is a reorganization of the perturbative series in $\alpha^{m-m^{*}}\frac{p_{m}(t)}{p_{m^{*}}(t)}\ll 1$ for each $m$. If the $p_{m}$ contain large logarithms, the RDF representation encodes them to all orders through the $g$-functions, but only those particular logarithmic structures implied by the fixed-order input, and should not be confused with the canonical NnLL logarithmic resummation structure in QCD derived from factorization and renormalization group (RG) methods Bauer:2000yr ; Banfi:2004yd . The choice of the higher-order component of $g$ effectively amounts to guessing this structure. It may be possible to extend the RDF to include this information (by e.g. imposing explicit $\mu$-dependence or factorization structure) in potential future work, but we do not pursue this here and content ourselves with fixed-scales and no assumed internal structure. This is resummation in the broad sense of formally reorganizing a perturbative series, of which the usual NnLL resummation is one specific type.
However, it is still highly tempting to draw a connection between the RDF and logarithmic resummation in QCD, especially since our ansatz consists of exponentiated polynomials in $\alpha$ and $\log(1/x)$. However, such a connection must be made with care and a few caveats. A logarithmic resummation for the CDF $P$ of an observable $x$ takes the form Catani:1992ua ; Banfi:2004yd  [^11] We have altered the notation and sign conventions of Ref. Catani:1992ua to better match our own. What they call $G_{nm}$”, we call $F_{mn}$” (note the swap between $m$ and $n$), since these coefficients are best associated with the integrals of our $f$-functions, $F=\int dt\,f$.:
@@
where $C_{m}$ and $F_{mn}$ are coefficients and $R$ is a remainder function that goes to 0 as $x\to 0$. A leading log (LL) calculation is one that includes all terms of the form $\alpha^{m}\log^{m+1}(1/x)$, a next-to-leading log (NLL) calculation is one that includes all terms of the form $\alpha^{m}\log^{m}(1/x)$, and so on to define NnLL. A related concept is the double log (DL) calculation, which includes the $\alpha\log^{2}(1/x)$ piece in the exponent at fixed coupling (unlike a true LL calculation which includes an entire tower of terms and the effects of running couplings). We can compare the formula for the logarithmic resummation of the CDF to the integral of the RDF, which is of the form (converting back to $x$-coordinates rather than $t$), which is given by
@@
However, extreme care must be taken in interpreting a RDF matching as an NnLL calculation. First, one must guarantee that all logarithms of the desired order are present in the fixed-order calculation — for example, at NLL, this includes effects due to the running of $\alpha_{s}$ and non-global logarithms, which are nontrivial to treat. Without this, there is no guarantee that matching to a fixed-order calculation will determine all the $F_{mn}$ needed to claim NnLL accuracy. Often, one only has the ability to compute a subset of the logarithms at any given order (see e.g. “Modified-Leading Logarithm (MLL) calculations” Dokshitzer:1992jv that miss some logarithms, or non-global logs Dasgupta:2001sh ), and we will see in Sec. 5.2 an explicit example where “leading order” does not imply “leading logs” in the case of simultaneous observables. Second, the non-matched part of $g$ controls all higher terms in the expansion. With no additional information or prior, this is effectively a random choice on the coefficients of the higher-order terms, which represents only a partial summation of all logarithms of that order. One must choose $g$ carefully to set these terms to zero if the goal is to compare to an NnLL calculation and look solely at the logarithms up to a given order.

### v2

@@ Abstract
Fixed-order perturbative calculations for differential cross sections can suffer from non-physical artifacts: they can be non-positive, non-normalizable, and non-finite, none of which occur in experimental measurements. We propose a framework, the Resummed Distribution Function (RDF), that, given a perturbative calculation for an observable to some finite order in $\alpha_{s}$, will “resum” the expression in a way that is guaranteed to match the original expression order-by-order and be positive, normalized, and finite. Moreover, our ansatz parameterizes all possible finite, positive, and normalized completions consistent with the original fixed-order expression, which can include NnLL resummed expressions. The RDF also enables a more direct notion of perturbative uncertainties, as we can directly vary higher-order parameters and treat them as nuisance parameters. We demonstrate the power of the RDF ansatz by matching to thrust to $\mathcal{O}(\alpha_{s}^{3})$ and extracting $\alpha_{s}$ with perturbative uncertainties by fitting the RDF to ALEPH data.
@@ 1 Introduction
However, at any finite order $M$, our physical perturbative predictions for $p(x)$ may violate unitarity by being non-normalizable or negative. Moreover, the perturbative expansion can be spoiled if $p_{m}(x)\sim\frac{1}{\alpha^{m}}$, which is especially the case in quantum chromodynamics (QCD) where large logarithms due to infrared and collinear divergences can arise. Additionally, while the perturbative series will be an integrable distribution, it will rarely be a proper function, and it may contain objects such as $\delta$-functions or ++-functions (which can technically restore unitarity or cancel divergences). Experimental measurements of $p(x|\alpha)$ are typically true functions that are finite everywhere as nature carries out all-orders calculations, and no experimental histogram will contain these objects.
@@
In this paper, we introduce a new ansatz, which we call the Resummed Distribution Function (RDF), for parameterizing proper PDFs of random variables. This ansatz is guaranteed to be positive, finite, and normalized, and thus satisfy unitarity. The RDF (which we denote as $q(x|\alpha)$ where $x$ can be either a single observable or a set of several observables) is an all-orders expansion in $\alpha$, and it effectively “resums” the fixed-order series using only unitarity and analyticity in $\alpha$ as consistency conditions. Given a fixed-order expansion for $p(x|\alpha)$ up to order $M$, our ansatz $q(x|\alpha)$ is guaranteed to match $p(x|\alpha)$ up to order $M$, meaning that it contains all of the information in the corresponding $p_{\rm FO}(x|\alpha)$. Through this ansatz, we effectively parameterize, in terms of free parameters, all possible higher-order extensions of $p(x|\alpha)$ permissible by unitarity. The RDF provides an automated approach for generating these potential “all-orders resummations” without any appeal to the structure of the theory, making it especially useful if a proper resummation is difficult or unavailable.
@@
The RDF provides an efficient machine-learning-inspired [^1] We say “inspired” because while the RDF is a universal probability estimator with parameters that can be numerically fit with gradient descent like many machine learning methods, our aim is to build a primarily analytical intuition for it. framework to learn normalized, finite, and conditional multivariable PDFs from analytic calculations, simulations, or real data, complementary to established machine-learning-and-adjacent methods for density estimation. We emphasize the strong inductive bias of the RDF ansatz: while it is a universal PDF approximator, it is particularly useful for distributions that admit a perturbative expansion in $\alpha$, as is often the case in a QFT. In contrast to normalizing flows [^2] We strongly considered naming our ansatz the “Re-normalizing flow”. papamakarios2021normalizingflowsprobabilisticmodeling and diffusion models ho2020denoisingdiffusionprobabilisticmodels , our ansatz is extremely lightweight: it can fit to numeric samples in a matter of minutes, requiring only a small set ($\sim\mathcal{O}(10)$) of trainable parameters. The RDF has a manifest expansion in $\alpha$ where terms of the expansion are directly parameterized, unlike normalizing flows and diffusion models where the expansion is not manifest or necessarily well-behaved. Compared to non-parametric methods for PDF estimation such as Kernel Density Estimation kde , our method can easily be made conditional on a parameter $\alpha$.
@@ 2 Resummed Distribution Functions: A Normalized, Positive, and Finite Ansatz for PDFs
Here, the functions $g$ can be any functions that (a) are bounded from above as any $t_{i}\to\infty$ and (b) are either analytic in $\alpha$ or have single-log dependence on $\alpha$. This parameterization ensures that the $f$ functions are always positive and analytic in $\alpha$, which is necessary for our ansatz to be a valid and analytic PDF (which we discuss further in the following section). [^4] Note that the functions $f$ and $g$ are ordinary functions of $t$ and $\alpha$, not PDFs of $t$ conditioned on $\alpha$. In Eq. (3), the $j$’th integral is only integrating over the $t_{j}$ argument of $f_{j}$, and all other arguments are not integrated over. The ansatz is completely specified by a choice of $g$ functions. Any choice of $g$ functions satisfying the above two conditions is a valid PDF, and all PDFs correspond to some choice of $g$.
@@ 2.1 Univariate Taylor-Expandable Probability Distributions
where $g_{mn}$ are coefficients, and $N$ is the highest power of $t$ considered. The only requirement is that $g_{mN}<0$ for normalization so that $g$ remains bounded from above. With this parameterization, each part $g$ is a series of polynomials in $t$ and $\alpha$, modulated by $\Theta$-functions in cases where $t$ does not necessarily start at zero. [^5] We draw a comparison here to Ref. Assi:2025ibi , which also involves determining parameters in an exponential to match to a calculation. However, Ref. Assi:2025ibi matches to already-resummed moments, and our $g$’s are inside of an exponential-integral-exponential structure rather than a plain single exponential.
@@ 2.2 Multivariate Taylor-Expandable Probability Distributions
The ordering of the $t_{i}$ is arbitrary, and thus this decomposition is not unique. [^6] Alternatively, one can choose to take an appropriately normalized sum over all possible orderings, but this is expensive computationally and does not offer any great advantage in our case.
@@ 2.3 Analytic Matching
The higher-order information in the ansatz will enforce normalization, positivity, and finiteness, effectively performing an “all-orders resummation”. For the following analytic matching procedure, we assume that we have access to an explicit expression for $p_{\rm FO}(t|\alpha)$ and the coefficient functions $p_{m}$ as defined in Eq. (1). Moreover, we require that the $p_{m}$ are indeed the Taylor coefficient functions of a well-defined distribution; that is, there exists an actual all-orders distribution the $p_{m}$ eventually converge to. [^7] This is not the case if, for example, one has $p_{0}(t)=1$ and $p_{m}(t)=0$ for all $m>0$. This does not converge to a valid distribution over $t$. Note that, in actual QFT calculations and especially QCD, this assumption may not be true: not only is there nonperturbative physics, but often perturbative expansions are only asymptotic series that do not actually converge. These issues are not unique to our method and we do not claim to solve either problem — the RDF is only as good as the perturbation theory used to derive $p_{\rm FO}$.
@@
Notice the $\mathcal{O}(\alpha^{M+1-m^{*}})$ inside the parentheses — since we have divided out $\alpha^{m^{*}}$, the expression inside the parentheses is an $\mathcal{O}(\alpha^{M-m^{*}})$ expression. We highlight this as it will be essential in our $\alpha$-counting. Substituting Eq. (16) into Eq. (15), we have
@@ 2.4 Numeric matching
where the first term corresponds to $\log(g^{*}(t,\alpha))$ and the second term to $g_{\rm Analytic}(t,\alpha)$. The numeric RDF differs from the analytic formula of Eq. (11) in a few minor ways. First, we allow the argument of the logarithm to be an arbitrary polynomial in $\alpha$, rather than using only the lowest power. We also force it to be strictly positive using with an absolute value function. [^9] Why absolute value and not anything else, e.g. a smoothed-out ReLU agarap2019deeplearningusingrectified to enforce positivity? We tried, and it was not as numerically stable. Second, we include explicit factorial scaling of the $g_{mn}$ coefficients for numerical stability. Lastly, we “smooth out” the $\Theta$-functions and absolute value functions, as indicated by the subscript $T$. We do this with the replacements $\Theta_{T}(x)\to\sigma(x;T)=\frac{1}{1+e^{-x/T}}$ and $|x|_{T}\to 2\sigma(\frac{x}{T})x-x$, with learnable parameters $T^{*}_{m}$, $T_{{\rm Analytic},m}$, and $T_{{\rm Abs},m}$ respectively per each $m$. This ensures that the derivatives of the RDF are well-defined at all points in $t$-space, which is especially important for the $\theta_{m}$’s to be learnable via gradient descent. All in all, we fit seven different objects: $g^{*}_{mn}$, ${g_{\mathrm{Analytic}}}_{mn}$, $\theta^{*}_{m}$, ${\theta_{\mathrm{Analytic}}}_{m}$, $T^{*}_{m}$, $T_{\mathrm{Analytic},m}$, and $T_{\textrm{Abs},m}$. For convenience, we will refer to this set of seven learnable objects as $\phi$.
@@
The binwise error in the denominator comes from uncertainties on the data or MC. We then backpropagate to calculate gradients of the loss with respect to the parameters and update the $g_{mn}$ and $\theta_{m}$ arrays with a gradient optimizer, in our case Adam adam .
@@ 3 What, When, and Why RDFs?
In this section, we hope to make clear what the RDF can and cannot do. It is important to be careful about the interpretation of the RDF, as it is not a method for obtaining extra information beyond unitarity for “free” or for performing automatic logarithmic resummation. At the lowest level, the only thing an RDF does is answer the question: “Given a fixed-order calculation, plus the additional knowledge that the all-orders distribution should be a proper probability distribution, what are all probability distributions consistent with that calculation?” The non-matched part of $g$ is infinitely flexible, and so without any additional information, further calculation, or a Bayesian prior, one cannot pick out which of these distributions is the correct one.
@@
However, there is some information gain due to the unitarity assumption: unitarity is an all-orders statement, so the RDF effectively provides a constraint on what the higher-order terms could be. This is indeed a nontrivial constraint: for example, after matching with an RDF to order $M$, and Taylor-expanding to see how the higher order $q_{m}$ for $m>M$ depend on the choice of $g$, one will never find that $q_{m}=0$ for all m>Mm>M [^10] Except in the trivial case where the original fixed-order expression happened to have no $\alpha$-dependence and was normalized to begin with., regardless of the choice of $g$, since that would violate unitarity. No matter what the original $p_{m}$ are, the higher-order coefficients must always conspire to make the entire distribution positive and normalized, which a generic all-orders guess may not do.
The RDF is an expansion in $\alpha^{m-m^{*}}\frac{p_{m}(t)}{p_{m^{*}}(t)}$, where $p_{m}(t)$ is the order-$m$ term in the $\alpha$ expansion and $p_{m^{*}}$ is the first nontrivial term in the expansion. Specifically, we assume in Eq. (18) that $\alpha^{m-m^{*}}\frac{p_{m}(t)}{p_{m^{*}}(t)}\ll 1$ in expanding the logarithms. This is the “resummation” in “Resummed Distribution Function”: it is a reorganization of the perturbative series in $\alpha^{m-m^{*}}\frac{p_{m}(t)}{p_{m^{*}}(t)}\ll 1$ for each $m$. If the $p_{m}$ contain large logarithms, the RDF representation encodes them to all orders through the $g$-functions, but only those particular logarithmic structures implied by the fixed-order input. This should not be confused with the canonical NnLL logarithmic resummation structure in QCD derived from factorization and renormalization group (RG) methods Bauer:2000yr ; Banfi:2004yd . The choice of the higher-order components of $g$ effectively amounts to guessing this structure. It may be possible to extend the RDF to include this information (by e.g. imposing explicit $\mu$-dependence or factorization structure) in potential future work, but we do not pursue this here and content ourselves with fixed-scales and no assumed internal structure. This is resummation in the broad sense of formally reorganizing a perturbative series, of which the usual NnLL resummation is one specific type.
However, it is still highly tempting to draw a connection between the RDF and logarithmic resummation in QCD, especially since our ansatz consists of exponentiated polynomials in $\alpha$ and $\log(1/x)$. Such a connection must be made with care and a few caveats. A logarithmic resummation for the CDF $P$ of an observable $x$ takes the form Catani:1992ua ; Banfi:2004yd  [^11] We have altered the notation and sign conventions of Ref. Catani:1992ua to better match our own. What they call $G_{nm}$”, we call $F_{mn}$” (note the swap between $m$ and $n$), since these coefficients are best associated with the integrals of our $f$-functions, $F=\int dt\,f$.:
@@
where $C_{m}$ and $F_{mn}$ are coefficients and $R$ is a remainder function that goes to 0 as $x\to 0$. A leading log (LL) calculation is one that includes all terms of the form $\alpha^{m}\log^{m+1}(1/x)$, a next-to-leading log (NLL) calculation is one that includes all terms of the form $\alpha^{m}\log^{m}(1/x)$, and so on to define NnLL. A related concept is the double log (DL) calculation, which includes the $\alpha\log^{2}(1/x)$ piece in the exponent at fixed coupling (unlike a true LL calculation which includes an entire tower of terms and the effects of running couplings). We can compare the formula for the logarithmic resummation of the CDF to the integral of the RDF, which is of the form (converting back to $x$-coordinates rather than $t$)
@@
However, extreme care must be taken in interpreting an RDF matching as an NnLL calculation. First, one must guarantee that all logarithms of the desired order are present in the fixed-order calculation — for example, at NLL, this includes effects due to the running of $\alpha_{s}$ and non-global logarithms, which are nontrivial to treat. Without this, there is no guarantee that matching to a fixed-order calculation will determine all the $F_{mn}$ needed to claim NnLL accuracy. Often, one only has the ability to compute a subset of the logarithms at any given order (see e.g. “Modified-Leading Logarithm (MLL) calculations” Dokshitzer:1992jv that miss some logarithms, or non-global logs Dasgupta:2001sh ), and we will see in Sec. 5.2 an explicit example where “leading order” does not imply “leading logs” in the case of simultaneous observables. Second, the non-matched part of $g$ controls all higher terms in the expansion. With no additional information or prior, this is effectively a random choice on the coefficients of the higher-order terms, which represents only a partial summation of all logarithms of that order. One must choose $g$ carefully to set these terms to zero if the goal is to compare to an NnLL calculation and look solely at the logarithms up to a given order.